A class of 195 students went on a field trip. They took 7 vehicles, some cars and some buses. Find the number of cars and the number of buses they took if each car holds 5 students and each bus hold 45 students?
Answer by angelmidnight(6)

x=cars
y=buses
x+y=7 vehicles
5x+45y= 195 students
-5x-5y=-35
+5x+45y=195
------------
40y=160
--- ---
40 40
y=4
x+4=7
x=3

They took 4 buses and 3 cars.